OpenStudy (anonymous):

Ms. Ellison and her boyfriend ate dinner at Chili's last night. Their total purchase was $27. When they received their bill , however, was included , and it was for $29.16. What was the tax rate?

OpenStudy (agreene):

27*x=29.16 solve for x

OpenStudy (anonymous):

??????

OpenStudy (agreene):

Cost of item * Tax = Total Cost $27 * x = $29.16 solve for x to find tax. x= 29.16/27 x= 27/25 x = 1.08